# AI on My Shoulder: Supporting Emotional Labor in Front-Office Roles with an LLM-based Empathetic Coworker

### v1

@@ 4.2.2. Linguistic Style & Semantics
To measure CDI, we computed the parts-of-speech of reframings using the Linguistic Inquiry and Word Count (LIWC) lexicon (Tausczik and Pennebaker, 2010). The CDI of Pro-Pilot’s messages was 281% more positive than that of Human’s (Cohen’s $d$=0.87)— indicating the language of Pro-Pilot was a lot more categorical. Zhou et al. found that when people are responding to personal incidents, such as bullying or venting, their support messages tend to have more negative CDI; in other words, they use dynamic language. However, they also found that in third-person reported events, such as news stories, support messages elicited more categorical language (Zhou et al., 2021). Pro-Pilot’s messages were more aligned to responses to reported events, but Human messages might be more effective if a CSR’s assessment of incivility is more personal.
Empathy.  Empathy refers to a cognitively complex process in which one can stand in the shoes of another person, to understand their perspective, emotions, and the situations they are in (Herlin and Visapää, 2016). Prior work evaluated the effectiveness of empathy in online interactions (Sharma et al., 2020) and chatbot interactions (Morris et al., 2018). Drawing on prior work, we employed a RoBERTa-based empathy detection model, finetuned on a dataset of empathetic reactions to news stories (Buechel et al., 2018; Tafreshi et al., 2021). Higher scores indicate a greater expression of empathy. Pro-Pilot’s messages scored 1.85% higher than Human’s — a small, but statistically significant effect (Cohen’s $d$=0.40). These results echo recent research on LLMs emulating empathy (Inzlicht et al., 2023; Kidder et al., 2024). Thus, Pro-Pilot shows promise in communicating support with an empathetic tone.
@@ 4.3. Psycholinguistic Analysis
Pronouns are indicative of interpersonal focus and narrative style. Pro-Pilot’s responses use significantly fewer first-person singular pronouns (I, us) (-67.36%, Cohen’s $d$=-0.58) and first-person plural pronouns (we, us) (-75.43%, Cohen’s $d$=-0.36), reflecting a less personal or collective identity focus. This reduction suggests that Pro-Pilot’s responses are less likely to include self-referential language, aligning with a more objective or detached communication style. These results further reinforce our results on dynamics (Section 4.2.2) that Pro-Pilot uses a relatively more objective but detached communication style.
@@ 4.4. Robustness
The results so far indicate that Pro-Pilot has promise in producing empathetic messages, but it begs the question: do we need a carefully crafted, domain-driven sequence of prompts (Section 3.2.2) for this? We replicated the analyses above with messages produced by zero-shot prompting of other LLMs— GPT-4 (gpt, 2024), GPT-4o (gpt, 2024), LLaMA-3.1 (lla, 2024), and Mistral-7B (Jiang et al., 2023). These LLMs vary in their architectures, training data, and optimization methods. We tested the differences using the Kruskal-Wallis test (McKight and Najab, 2010) and found Pro-Pilot’s responses to be closer to humans, more empathetic, and more controllable. The results of this benchmarking are reported in Table A1. Now, one might ask if Pro-Pilot’s messages, with all its linguistic distinctions, matter to the CSRs? To answer this, we checked the differences in CSR’s evaluation of Pro-Pilot and Human’s messages.
Pro-Pilot’s Perceived Empathy. Empathy is a complex phenomenon with many dimensions. To measure perceived empathy, we combined scales from previous studies (Sharma et al., 2023; Nambisan, 2011). In general, Pro-Pilot messages were perceived to be significantly more empathetic (Cohen’s $d$=0.35). Moreover, we also calculated differences in the subscales, after correcting for Bonferroni multiple pairwise-comparisons. In terms of raw averages, Pro-Pilot scored highest on sincerity and actionability. Messages from Pro-Pilot were considered more genuine and less pretentious (sincerity: Cohen’s $d$=0.31). CSRs also felt they could take practical action based on the messages (actionability: Cohen’s $d$=0.48).
5. Pro-Pilot: User Evaluation
After establishing that Pro-Pilot can produce situationally appropriate messages, we set out to study how CSRs might actually interact with such an AI-assistant. The critical nature of front-office work raised practical and ethical challenges of deploying a prototype into actual workflows. Therefore, to answer RQ II, we chose to design a simulation exercise where real CSRs could interact with uncivil clients using Pro-Pilot. We deployed Pro-Pilot as a technology probe (Hutchinson et al., 2003) — a functional piece of technology that is presented to learn its use and about the users.
@@ 5.1. Task Environment
We wanted to study the role of Pro-Pilot’s core component Emo-Reframe (Section 3) in realistic cases of incivility. Thus, we built a web-based prototype environment that resembles typical client interaction interfaces that a CSR might use. Additionally, to isolate the role of Emo-Reframe, we added two new components to Pro-Pilot that represent other forms of intelligent assistance:
Figure 4. Task interface for the user evaluation. The client names were randomly assigned.
@@
    Client Simulant: LLMs have been successfully applied in previous research to simulate challenging interlocutors (Shaikh et al., 2023). Following the same principle, we devised a simulation exercise where participants needed to interact with complaining clients. We reused the Client-Agent described in Section 3.1.2 to continue a conversation until their complaint is resolved, or 10-12 turns, whichever occurs sooner.
@@
    Info-Guide Panel: Typically, AI assistants at work provide problem solving support (Mozannar et al., 2024). We replicated this functionality by augmenting Pro-Pilot with LLM-powered generated troubleshooting guidelines. These were a trace of suggestions to help the CSR solve the specific complaint.
@@
    Emo-Label Panel: It is common for CSR interfaces to have sentiment tags highlighting how the client feels and anticipate their behaviors. A CSR’s ability to understand the emotional perspective of others can mitigate the negative effects of incivility  (Rafaeli et al., 2012; Rupp et al., 2008; Schutte et al., 2001). We implemented an ensemble sentiment classifier that uses soft-voting to combine the estimates from NLTK (Hardeniya et al., 2016), TextBlob (Loria et al., 2018), and Transformers (Wolf et al., 2020). The output was in the form of a 7-point scale, ranging from very negative to very positive.
@@
All of Pro-Pilot’s panels were placed on the sides to keep them glanceable, while the main chat remained the primary area of focus. These panels are dynamically updated after every new response from the client. For the purposes of the study, we included a simple semantic differential Likert scale under each panel to ensure that the participant reads them carefully. Pro-Pilot also provided participants and response cues (short phrases) to nudge replies to the client. Fig. 4 provides an overview of the environment with all its components.
@@ 5.2. Study Protocol
Every user session was facilitated over Zoom and lasted at most 90 minutes. All interviews were led by the first author, with the second author observing. Participants accessed the simulation exercise via a unique URL to our study portal. They were tasked to role-play as a CSR who needs to resolve client complaints. The session was divided into two phases. First, participants completed the simulation exercise. To emulate realistic workplace demands, participants were informed that the clients would be rating their complaint resolution skills, and this subjective rating would determine their bonus compensation. Then, they reflected on their interaction with Pro-Pilot through our interview. The bonus incentive was minor deception, and all participants were paid the full \$$\$50$ in the form of a gift card.
@@
5.2.1. Simulation Exercise
In line with prior studies of front-office worker behavior (Frey-Cordes et al., 2020), each participant was randomly assigned a domain, either airlines or hotels. Before starting the exercise, participants completed a pre-task survey describing their experience interacting with their typical client. We included multiple instruments to capture how the client treated them (Spencer and Rupp, 2009), the cognitive demands and resources available to them (Demerouti et al., 2001), and how they affectively perceived the conversations (Betella and Verschure, 2016). To familiarize participants with the interface, they first interacted with a civil client with only Info-Guide supporting them. Once they became accustomed to this process, they proceeded to the main exercise. Participants needed to handle complaints from three clients — one civil and two uncivil. They only received suggestions and insight from Emo-Label, and Emo-Reframe for the last uncivil client. At the end of every client interaction, they responded to the same survey questions as the pre-task survey. Additionally, they also reported how they perceived Pro-Pilot based on dimensions of AI-mediated support (Liu and Sundar, 2018). The survey measurements can be found in Section A.2. This iterative task setup let us compare participants’ experiences when dealing with uncivil clients, with and without emotional support, and to examine how these factors influenced their performance and emotional responses.
5.2.2. Semi-Structured Interview
After interacting with clients, participants proceeded with a semi-structured interview. They answered a series of open-ended questions to provide deeper insight into their primary goals when dealing with uncivil clients, their attitude towards emotional well-being, and their evaluation of Pro-Pilot, especially, in comparison to human coworker support.
5.3. Thematic Analysis
We performed inductive coding to identify Pro-Pilot’s role in front-office work (Braun and Clarke, 2006). Two authors carefully read each transcript and performed open-coding. The authors first coded the transcripts independently and then met frequently to reconcile disagreements. They iteratively improved the codes in 20% chunks. Next, we conducted affinity mapping of 258 initial codes using Miro. We merged similar codes together and pruned out codes outside the scope of our research. During clustering, we first organized concepts related to the socio-organizational norms of front-office work. Based on the gaps in these norms we clustered codes related to the capabilities of Pro-Pilot. Our main findings comprise 128 codes which were organized into a four-level thematic structure. We primarily elaborate on the broad themes of Utility, Complementarity, and Pitfalls by anchoring them in the normative patterns that emerged from Reactions to Client Incivility and Role of Coworkers in ER. Fig. 5 provides an overview of the main themes, and their relationship, that we cover in the remaining findings.
Figure 5. The user evaluation findings sections correspond to 3 major top-level themes describing Pro-Pilot. These themes were grounded in the normative patterns of client incivility and coworker support that were raised by our participants.
@@ 5.3. Thematic Analysis
Summary of user evaluation findings for Pro-Pilot’s role in supporting CSRs through client incivility and coworker support challenges.
6. Findings: Clarifying the Role of Pro-Pilot in Uncivil Interactions
The negativity of clients can be contagious (Barsade et al., 2018), and participants described their attitude turning negative (P06), lowered productivity (P08), feeling drained or depressed (P18, P20), and even wanting to reciprocate the incivility (P19). Furthermore, participants also recognized that client incivility can be a form of micro-trauma (Straussner and Calnan, 2014) that was only apparent after several shifts or even years (P14). Many of them work in the fear that a dissatisfied client might report them (P01) and subsequently feel trapped in their role (P04). Given these experiences, the participants in our study were uniquely positioned to assess Pro-Pilot.
Figure 6. Participants rated the cognitive demands of interacting with the uncivil Client-Agent lower and similar to civil Client-Agent when they were assisted by Pro-Pilot’s Emo-Reframe
@@ 6. Findings: Clarifying the Role of Pro-Pilot in Uncivil Interactions
Density plot showing that participants rated the cognitive demands of interacting with the uncivil Client-Agent lower and similar to civil interactions when assisted by Pro-Pilot’s Emo-Reframe.
Figure 7. Participants rated the helpfulness of Pro-Pilot after every message; Emo-Reframe was significantly better than Emo-Label and similar to Info-Guide (crossing zero-line).
[…]
Mean pairwise difference plot showing that participants rated Pro-Pilot’s Emo-Reframe as significantly more helpful than Emo-Label and comparable to Info-Guide.
Relatability of the Simulation: Participants had to interact with an uncivil client that replicates emotionally challenging situations of front-office work. During the interactions, several participants exhibited observable reactions to incivility, such as defensively laughing (P06, P18), eye-rolling (P10), and verbally labeling the client (e.g., “spicy”–P05). P07 even claimed his “adrenaline was peaking.” P16 endorsed our simulation by stating, “it was a little intense, but that’s definitely how people react sometimes; so, it felt very real.”

### v2

@@ 4.2.2. Linguistic Style & Semantics
To measure CDI, we computed the parts-of-speech of reframing messages using the Linguistic Inquiry and Word Count (LIWC) lexicon (Tausczik and Pennebaker, 2010). The CDI of Pro-Pilot’s messages was 281% more positive than that of Human’s (Cohen’s $d$=0.87)— indicating the language of Care-Pilot was a lot more categorical. Zhou et al. found that when people are responding to personal incidents, such as bullying or venting, their support messages tend to have more negative CDI; in other words, they use dynamic language. However, they also found that in third-person reported events, such as news stories, support messages elicited more categorical language (Zhou et al., 2021). Care-Pilot’s messages were more aligned to responses to reported events, but Human messages might be more effective if a CSR’s assessment of incivility is more personal.
Empathy.  Empathy refers to a cognitively complex process in which one can stand in the shoes of another person, to understand their perspective, emotions, and the situations they are in (Herlin and Visapää, 2016). Prior work evaluated the effectiveness of empathy in online interactions (Sharma et al., 2020) and chatbot interactions (Morris et al., 2018). Drawing on prior work, we employed a RoBERTa-based empathy detection model, fine-tuned on a dataset of empathetic reactions to news stories (Buechel et al., 2018; Tafreshi et al., 2021). Higher scores indicate a greater expression of empathy. Care-Pilot’s messages scored 1.85% higher than Human’s — a small, but statistically significant effect (Cohen’s $d$=0.40). Since empathy is a core mechanic for Care-Pilot, we further tested it with another RoBERTa-based empathy classifier, which was trained on a dataset of mental health peer-support (Sharma et al., 2020). A CSR is unable to explicitly seek support while also attending to a client, thus, for this classifier we used the output of the thought subsequence (Section 3.2.2) as a proxy for their support seeking message. We inspected the two sets empathetic response messages for expressions of “emotions such as warmth, compassion, and concern”, or as Sharma et al. describe it, emotional reactivity  [^6] Sharma et al.’s classifier can also label interpretations and explorations, as these are relevant to peer support (Sharma et al., 2020). Our scenario differs as users (CSRs) do not have back-and-forth communication with Care-Pilot or Human.. Care-Pilot’s messages had a significantly higher emotional reactivity than Human’s (12.52%; Cohen’s $d$=0.31). On closer look, we found that Human messages were often classified to contain no emotional reactivity, but also included instances of strong emotional reactivity. In contrast, Care-Pilot messages were consistently classified to contain weak emotional reactivity which was equivalent to a score of $1.00$. For further context, Sharma et al.’s findings show that emotional reactivity in peer support groups on Reddit varied from $0.70-0.45$. For participants in our dataset the average score was $0.88$. These results echo recent research on LLMs emulating empathy (Inzlicht et al., 2023; Kidder et al., 2024; Saha et al., 2025). Thus, Care-Pilot shows promise in communicating support with an empathetic tone.
@@ 4.3. Findings: Psycholinguistic Analysis
Pronouns are indicative of interpersonal focus and narrative style. Pro-Pilot’s responses use significantly fewer first-person singular pronouns (I, us) (-67.36%, Cohen’s $d$=-0.58) and first-person plural pronouns (we, us) (-75.43%, Cohen’s $d$=-0.36), reflecting a less personal or collective identity focus. This reduction suggests that Pro-Pilot’s responses are less likely to include self-referential language, aligning with a more objective or detached communication style. These results further reinforce our results on dynamics (Section 4.2.2) that Care-Pilot uses a relatively more objective but detached communication style.
@@ 4.4. Robustness
The results so far indicate that Care-Pilot has promise in producing empathetic messages, but it begs the question: do we need a carefully crafted, domain-driven sequence of prompts (Section 3.2.2) for this? We replicated the analyses above with messages produced by zero-shot prompting of other LLMs— GPT-4 (gpt, 2024), GPT-4o (gpt, 2024), LLaMA-3.1 (lla, 2024), and Mistral-7B (Jiang et al., 2023). These LLMs vary in their architectures, training data, and optimization methods. We tested the differences using the Kruskal-Wallis test (McKight and Najab, 2010) and found Care-Pilot’s scores across metrics to be closer to humans, more empathetic, and more controllable. The results of this benchmarking are reported in Table D1. Specifically, zero-shot prompting often led to greater verbosity and variation in messages, whereas Care-Pilot offered a more deterministic solution.
Care-Pilot’s Perceived Empathy. Now, one might ask if Care-Pilot’s messages, with all its linguistic distinctions, matter to the CSRs? To answer this, we checked the differences in CSR’s evaluation of Care-Pilot and Human’s messages. Empathy is a complex phenomenon with many dimensions. To measure perceived empathy, we combined scales from previous studies (Sharma et al., 2023; Nambisan, 2011). In general, Care-Pilot messages were perceived to be significantly more empathetic (Cohen’s $d$=0.35). Moreover, we also calculated differences in the subscales, after correcting for Bonferroni multiple pairwise-comparisons. In terms of raw averages, Care-Pilot scored the highest on sincerity and actionability. Messages from Care-Pilot were considered more genuine and less pretentious (sincerity: Cohen’s $d$=0.31). CSRs also felt they could take practical action based on the messages (actionability: Cohen’s $d$=0.48).
5. Care-Pilot: User Evaluation
After establishing that Care-Pilot can produce situationally appropriate messages, we set out to study how CSRs might actually interact with such an AI-assistant. The critical nature of front-office work raised practical and ethical challenges of deploying a prototype into actual workflows. Therefore, to answer RQ II, we chose to design a simulation exercise where real CSRs could interact with uncivil clients using Care-Pilot. We deployed Care-Pilot as a technology probe (Hutchinson et al., 2003) — a functional piece of technology that is presented to learn its use and about the users.
@@ 5.1. Task Environment
We wanted to study the role of Care-Pilot’s core component Emo-Reframe (Section 3) in realistic cases of incivility. Thus, we built a web-based prototype environment that resembles typical client interaction interfaces that a CSR might use. Additionally, to isolate the role of Emo-Reframe, we added two new components to Care-Pilot that represent other forms of intelligent assistance:
Figure 4. Task interface for the user evaluation. The client names were randomly assigned. Appendix E.1 contains dedicated figures of the major components of this interface for easier reading.
@@
    Client Simulant: LLMs have been successfully applied in previous research to simulate challenging interlocutors (Shaikh et al., 2023). Following the same principle, we devised a simulation exercise where participants needed to interact with complaining clients. We reused the Client-Agent described in Section 3.1.2 to continue a conversation until their complaint is resolved, or 10–12 turns, whichever occurs sooner.
@@
    Info-Guide Panel: Typically, AI assistants at work provide problem-solving support (Mozannar et al., 2024). We replicated this functionality by augmenting Care-Pilot with LLM-generated troubleshooting guidelines. These were a trace of suggestions to help the CSR solve the specific complaint.
@@
    Emo-Label Panel: It is common for CSR interfaces to have sentiment tags highlighting how the client feels and anticipate their behaviors. A CSR’s ability to understand the emotional perspective of others can mitigate the negative effects of incivility  (Rafaeli et al., 2012; Rupp et al., 2008; Schutte et al., 2001). In fact, service management research has started promoting emotion or sentiment recognition features for CSRs (Henkel et al., 2020). We implemented an ensemble sentiment classifier that uses soft-voting to combine the estimates from NLTK (Hardeniya et al., 2016), TextBlob (Loria et al., 2018), and Transformers (Wolf et al., 2020). The output was in the form of a 7-point scale, ranging from very negative to very positive.
@@
All of Care-Pilot’s panels were placed on the sides to keep them glanceable, while the main chat remained the primary area of focus. These panels are dynamically updated after every new response from the Client-Agent. For the purposes of the study, we included a simple semantic differential Likert scale under each panel to measure helpfulness of support (further details in Appendix E.2.1). Requiring paritipants to respond to this scale before continuing the conversation ensured that they read the support messages carefully. Care-Pilot also provided participants and response cues (short phrases) to nudge replies to the client. Fig. 4 provides an overview of the environment with all its components.
@@ 5.2. Study Protocol
Every user session was facilitated over Zoom and lasted at most 90 minutes. All interviews were led by the first author, with the second author observing. Participants accessed the simulation exercise via an online portal. They were tasked to role-play as a CSR who needs to resolve client complaints. The session was divided into two phases. First, participants completed the simulation exercise. To emulate realistic workplace demands, participants were informed that the clients would be rating their complaint resolution skills, and this subjective rating would determine their bonus compensation. Then, they reflected on their interaction with Care-Pilot through our interview. The bonus incentive was minor deception, and all participants were paid the full \$$\$50$ in the form of a gift card.
@@
Summary of participant demographic for user evaluation. ID Gender Age Race Work Sector P01 Female 21-29 Asian P02 Female 50-59 White Online Business P03 Male 30-39 Black or AA Real estate P04 Male 30-39 Asian Healthcare P05 Female 30-39 Black or AA Insurance P06 Female 30-39 White Transportation, Logistics P07 Male 30-39 White Education P08 Female 40-49 Asian Finance P09 Male $\geq$60 White Education, Airlines P10 Female 30-39 Asian Consulting P11 Male 21-29 Asian Data Science P12 Male 21-29 Asian Finance P13 Female 21-29 White Government P14 Female 40-49 White Government P15 Male 21-29 Black or AA Technology P16 Female 30-39 White Accounting, Logistics P17 Male 40-49 Black or AA Sports P18 Female 40-49 White Finance P19 Female 18-20 Black or AA Sales P20 Male 30-39 Black or AA Real estate
5.2.1. Simulation Exercise
In line with prior studies of front-office worker behavior (Frey-Cordes et al., 2020), each participant was randomly assigned a domain, either airlines or hotels. Before starting the exercise, participants completed a pre-task survey describing their experience interacting with their typical client. We included multiple instruments to capture how the client treated them (Spencer and Rupp, 2009), the cognitive demands and resources available to them (Demerouti et al., 2001), and how they affectively perceived the conversations (Betella and Verschure, 2016). To familiarize participants with the interface, they first interacted with a civil client with only Info-Guide supporting them. Once they became accustomed to this process, they proceeded to the main exercise. Participants needed to handle complaints from three clients — one civil and two uncivil. They only received suggestions and insight from Emo-Label, and Emo-Reframe for the last uncivil client. If we count 1 turn as an exchange where the Client-Agent sends a message and receives a response from the CSR, then the average number of turns while conversing with civil clients was $4.65$ and lasted $9.22$ minutes. Conversing with uncivil Client-Agent with only Info-Guide for assistance took longer as they lasted an average of $6.16$ turns and $12.11$ minutes. When Emo-Reframe was available to assist, then conversations with uncivil Client-Agent took $5$ turns and $12.16$ minutes on average. At the end of every client interaction, they responded to the same survey questions as the pre-task survey. Additionally, they also reported how they perceived Care-Pilot based on dimensions of AI-mediated support (Liu and Sundar, 2018). The survey measurements can be found in Appendix E.2. This iterative task setup let us compare participants’ experiences when dealing with uncivil clients, with and without emotional support, and to examine how these factors influenced their performance and emotional responses.
5.2.2. Semi-Structured Interview
After interacting with clients, participants proceeded with a semi-structured interview. They answered a series of open-ended questions to provide deeper insight into their primary goals when dealing with uncivil clients, their attitude towards emotional well-being, and their evaluation of Care-Pilot, especially, in comparison to human coworker support.
5.3. Thematic Analysis
Figure 5. The user evaluation findings sections correspond to 3 major top-level themes describing Care-Pilot. These themes were grounded in the normative patterns of client incivility and coworker support that were raised by our participants.
@@ 5.3. Thematic Analysis
Summary of user evaluation findings for Care-Pilot ’s role in supporting CSRs through client incivility and coworker support challenges.
We performed inductive coding to identify Care-Pilot’s role in front-office work (Braun and Clarke, 2006). Two authors carefully read each transcript and performed open-coding. The authors first coded the transcripts independently and then met frequently to reconcile disagreements. They iteratively improved the codes in 20% chunks. Next, we conducted affinity mapping of 258 initial codes using Miro. We merged similar codes together and pruned out codes outside the scope of our research. During clustering, we first organized concepts related to the socio-organizational norms of front-office work. Based on the gaps in these norms, we clustered codes related to the capabilities of Care-Pilot. Our main findings comprise 128 codes, which were organized into a four-level thematic structure. We primarily elaborate on the broad themes of Utility, Complementarity, and Pitfalls by anchoring them in the normative patterns that emerged from Reactions to Client Incivility and Role of Coworkers in ER. Fig. 5 provides an overview of the main themes, and their relationship, that we cover in the remaining findings.
6. Findings: Clarifying the Role of Care-Pilot in Uncivil Interactions
The negativity of clients can be contagious (Barsade et al., 2018), and participants described their attitude turning negative (P06), lowered productivity (P08), feeling drained or depressed (P18, P20), and even wanting to reciprocate the incivility (P19). Furthermore, participants also recognized that client incivility can be a form of microtrauma (Straussner and Calnan, 2014) that was only apparent after several shifts or even years (P14). Many of them work in the fear that a dissatisfied client might report them (P01) and subsequently feel trapped in their role (P04). Given these experiences, the participants in our study were uniquely positioned to assess Care-Pilot.
Figure 6. Participants rated the cognitive demands of interacting with the uncivil Client-Agent lower and similar to civil Client-Agent when they were assisted by Care-Pilot’s Emo-Reframe
@@ 6. Findings: Clarifying the Role of Care-Pilot in Uncivil Interactions
Density plot showing that participants rated the cognitive demands of interacting with the uncivil Client-Agent lower and similar to civil interactions when assisted by Care-Pilot’s Emo-Reframe.
Figure 7. Participants rated the helpfulness of Care-Pilot after every message; Emo-Reframe was significantly better than Emo-Label and similar to Info-Guide (crossing zero-line).
@@
Mean pairwise difference plot showing that participants rated Care-Pilot ’s Emo-Reframe as significantly more helpful than Emo-Label and comparable to Info-Guide.
Relatability of the Simulation: Participants had to interact with an uncivil client that replicates uncivil situations in front-office work. During the interactions, several participants exhibited observable reactions to incivility, such as defensively laughing (P06, P18), eye-rolling (P10), and verbally labeling the client (e.g., “spicy”–P05). P07 even claimed his “adrenaline was peaking.” P16 endorsed our simulation by stating, “it was a little intense, but that’s definitely how people react sometimes; so, it felt very real.”